Bramble Corporation issues $400,000 of 11% bonds that are due in 9 years and pay interest semi-annually. At the time of issue, the market rate for such bonds is 10%.

Click here to view the factor table PRESENT VALUE OF 1.
Click here to view the factor table PRESENT VALUE OF AN ANNUITY OF 1.

Calculate the bonds’ issue price by using (1) factor Tables A.2 and A.4, (2) a financial calculator, or (3) Excel function PV. (Hint: Refer to Chapter 3 for tips on calculating.) (For calculation purposes, use 5 decimal places as displayed in the factor table provided and round final answer to 0 decimal places, e.g. 5,275.)

Issue price of bond $enter the Issue price of bond in dollars rounded to 0 decimal places
Answer #1
Issue price of bond 423379
Workings:
Semi-annual interest 22000 =400000*11%*6/12
Semi-annual interest rate 5% =10%/2
Number of periods 18 =9*2
Amount PV factor 5% Present value
Semi-annual interest 22000 11.68959 257171
Principal 400000 0.41552 166208
Total 423379
PV factor 5%
Semi-annual interest 11.68959 =(1-(1.05)^-18)/0.05
Principal 0.41552 =1/1.05^18